# Q-Insight: Understanding Image Quality via Visual Reinforcement Learning

### v1

@@ Abstract
Image quality assessment (IQA) focuses on the perceptual visual quality of images, playing a crucial role in downstream tasks such as image reconstruction, compression, and generation. The rapid advancement of multi-modal large language models (MLLMs) has significantly broadened the scope of IQA, moving toward comprehensive image quality understanding that incorporates content analysis, degradation perception, and comparison reasoning beyond mere numerical scoring. Previous MLLM-based methods typically either generate numerical scores lacking interpretability or heavily rely on supervised fine-tuning (SFT) using large-scale annotated datasets to provide descriptive assessments, limiting their flexibility and applicability. In this paper, we propose Q-Insight, a reinforcement learning-based model built upon group relative policy optimization (GRPO), which demonstrates strong visual reasoning capability for image quality understanding while requiring only a limited amount of rating scores and degradation labels. By jointly optimizing score regression and degradation perception tasks with carefully designed reward functions, our approach effectively exploits their mutual benefits for enhanced performance. Extensive experiments demonstrate that Q-Insight substantially outperforms existing state-of-the-art methods in both score regression and degradation perception tasks, while exhibiting impressive zero-shot generalization to comparison reasoning tasks. Code will be available at https://github.com/lwq20020127/Q-Insight.
@@ 1 Introduction
Figure 1: We present the PLCC comparison between proposed Q-Insight and other IQA metrics on the left and list the applications of our Q-Insight on the right. Q-Insight shows significant performance on the out-of-domain datasets (e.g. CSIQ) compared to methods like DeQA, while simultaneously achieving score regression, degradation perception, and zero-shot image comparison reasoning tasks.
@@
Specifically, we design Q-Insight upon the GRPO framework. In our Q-Insight, we jointly optimize score regression and degradation perception tasks, and carefully design three reward functions: a verifiable score reward for the score regression task, and degradation classification and intensity perception rewards for the degradation perception task. Consequently, Q-Insight effectively exhibits robust reasoning performance in image quality assessment tasks using only limited Mean Opinion Scores (MOS) and degradation labels. As shown in Fig. 1, our Q-Insight shows exceptional improvements on OOD datasets, while demonstrating comprehensive quality perception and sophisticated reasoning abilities, such as recognizing that a slightly out-of-focus background, which might conventionally be perceived as undesirable, actually helps emphasize the main subject of the image. Our empirical investigation reveals that: (1) training solely with score labels results in poor perception of image detail degradations (e.g., JPEG compression), while jointly training with the degradation perception task significantly enhances the model’s sensitivity to such degradations, and (2) score regression and degradation perception tasks are mutually beneficial. Extensive experiments across score regression and degradation perception tasks demonstrate that Q-Insight consistently outperforms existing model-based IQA metrics as well as SFT-driven large language models. Moreover, it exhibits impressive zero-shot generalization on unseen tasks, such as image comparison reasoning, highlighting the robustness and versatility of our method. In summary, our contributions are:
❑ (1) We propose Q-Insight, the first reasoning-style model specifically designed for comprehensive image quality understanding. Unlike previous methods that depend heavily on detailed textual descriptions for supervised fine-tuning (SFT), our approach achieves superior understanding capability using only limited mean opinion scores or degradation labels.
@@ 2 Related Work
Score-based IQA methods include full-reference and no-reference approaches. Full-reference methods [47, 39, 63] assess image quality by comparing distorted images with high-quality references using traditional metrics (e.g., SSIM [47]) or advanced deep-learning-based metrics [3, 4, 9, 8, 12, 34] like LPIPS [64]. Non-reference methods evaluate quality without reference images, shifting from traditional handcrafted statistics [26, 27, 28, 29, 30, 36] to deep-learning-derived quality priors [17, 18, 23, 32, 41, 71, 72, 42, 46]. Recent multi-modal large language model (MLLM)-based methods, such as Q-Align [51] and DeQA-Score [56], leverage MLLMs’ knowledge and perceptual abilities to produce accurate scores. However, they sacrifice the intrinsic descriptive capabilities of MLLMs.
@@ 3.1 Preliminaries
Group Relative Policy Optimization (GRPO) is an innovative reinforcement learning paradigm that has been widely used in models such as DeepSeek R1-Zero, achieving excellent results. Unlike PPO, which requires an explicit critic model to evaluate the performance of the policy model, GRPO directly computes the advantage by comparing a group of responses sampled from the policy model, greatly reducing the computational burden. Specifically, given a query $q$, GRPO samples $N$ distinct responses $\{o^{(1)},o^{(2)},\ldots,o^{(N)}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. Then, the method performs the corresponding actions and receives the respective rewards $\{r^{(1)},r^{(2)},...,r^{(N)}\}$ according to the task-specific rules. By calculating the mean and standard deviation of the rewards, the relative advantages of each response can be obtained as follows:
@@
where $\hat{A}^{(i)}$ represents the normalized relative quality of the $i$-th answer. Overall, GRPO guides the policy model to prioritize higher-quality answers that receive higher reward values within the group.
After obtaining the advantage $\hat{A}^{(i)}$, GRPO calculates the ratio of the probabilities of each response under the new policy $\pi_{\theta_{\text{new}}}$ and the old policy $\pi_{\theta_{\text{old}}}$, denoted as $\text{ratio}^{(i)}$. To prevent overly large updates to the model and stabilize training, GRPO restricts the $\text{ratio}^{(i)}$ to the range $[1-\delta,1+\delta]$. To further maintain closeness to the reference distribution $\pi_{\text{ref}}$, a KL divergence penalty weighted by $\beta$ is adopted. Finally, the optimization objective of GRPO can be formulated as follows:
$\mathcal{J}(\theta)=\mathbb{E}_{[q\sim Q,o^{(i)}\sim\pi_{\theta_{\text{old}}}]}$$\left\{\min\left[\text{ratio}^{(i)},\operatorname{clip}\left(\text{ratio}^{(i)},1-\delta,1+\delta\right)\right]\hat{A}^{(i)}-\beta\cdot\mathbb{D}_{\text{KL}}[\pi_{\theta_{\text{new}}}||\pi_{\text{ref}}]\right\},$
$with~{}~{}~{}\text{ratio}^{(i)}=\pi_{\theta_{\text{new}}}(o^{(i)}~{}|~{}q)~{}/~{}{\pi_{\theta_{\text{old}}}(o^{(i)}~{}|~{}q)},$ (2)
where $Q$ denotes the candidate question set, $\mathbb{D}_{\text{KL}}$ denotes the KL regularization. $\pi_{\text{ref}}$ is typically a frozen pre-trained M-LLM. In a nutshell, GRPO effectively integrates consistent policy updates and strong reward signals in a balanced and controlled way.
@@ 3.2 Overview of Q-Insight
Figure 2: Overview of the proposed Q-Insight framework. The policy model receives queries from multiple tasks and generates corresponding groups of responses accompanied by explicit reasoning steps. Task-specific reward functions ($R_{\text{scr}}$, $R_{\text{deg}}$, and $R_{\text{lev}}$) are then computed, and the policy model is subsequently optimized jointly using the multi-task group relative policy optimization algorithm.
@@ 3.3 Multi-Task Group Relative Policy Optimization
Reward design is crucial for RL-based approaches, as it allows Q-Insight to iteratively enhance its reasoning capabilities. Specifically, given each input data pair, the policy model $\pi_{\theta}$ generates a group of $N$ responses, denoted as $\{o^{(i)}\}_{i=1}^{N}$. We then evaluate these response groups using our designed reward functions to obtain $\{r^{(i)}\}_{i=1}^{N}$. This includes a general format reward shared by all tasks, along with task-specific rewards customized to the distinct characteristics of each task.
Format reward evaluates whether the reasoning steps are properly enclosed within “<think>” and “</think>” tags, and the final answer is correctly enclosed within “<answer>” and “</answer>” tags. Additionally, the content inside “<answer>” must follow a JSON-like format, beginning with “{”, ending with “}”, and containing no additional “{” or “}” characters internally. Finally, a reward score $r^{(i)}_{\text{fmt}}$ for $i$-th response of $1$ is assigned if all these conditions are satisfied. Otherwise, the reward is 00.
Rewards for score regression task. A standard way to quantify image quality is using the Mean Opinion Score (MOS). Inspired by DeepSeek-R1 [14], which leverages the final answers of math problems as a guide to promote deeper reasoning, we similarly employ MOS scores as general targets to encourage detailed reasoning and critical thinking when assessing image quality. Denoting the predicted score of the $i$-th response as $\text{scr}_{\text{pred}}^{(i)}$ and the ground-truth score as $\text{scr}_{\text{gt}}$, we design a verifiable score reward function $R_{\text{scr}}$ as follows. The reward value $r_{\text{scr}}^{(i)}$ for the $i$-th response is determined by:
@@
where $\epsilon$ is a predefined threshold, allowing the model’s score predictions to fluctuate within an acceptable range rather than requiring exact accuracy. Thus, the predicted score receives a reward of $1$ if it falls within the threshold $\epsilon$ of the ground-truth MOS, and 00 otherwise.
Table 1: Prompts for Different Tasks. The system prompt is shared across all tasks, while task-specific prompts are additionally designed for each individual task.
Prompt for Reference-Based Zero-shot Comparison Reasoning: Given reference low-quality image: <ref_image>, Image A: <image_A> and Image B: <image_B>. Image A and Image B are results from two resolution methods. Explain and justify whether Image A or Image B is better. You should think from the following aspects: (1) Content fidelity and consistency with the reference image and (2) overall quality. Your answer should be "Image A" or "Image B".
Rewards for degradation perception task. In the degradation perception task, the model is expected to predict both the distortion class and its corresponding level. Denoting the predicted distortion class and level of the $i$-th response as $\text{deg}^{(i)}_{\text{pred}}$ and $\text{lev}^{(i)}_{\text{pred}}$ respectively, we define the degradation classification reward as follows. The reward value $r^{(i)}_{\text{deg}}$ is determined by:
@@
This means the predicted distortion class receives a reward of $1$ if correct, and 00 otherwise. Similarly, the distortion-level reward $r^{(i)}_{\text{lev}}$ is determined by:
@@
Thus, the predicted distortion level earns a reward of $1$ only if both the predicted class and the level exactly match the ground truth; otherwise, it receives 00.
Finally, the overall reward of $i$-th response is calculated as:
@@
where $\mathds{1}_{\text{scr}}$ equals $1$ if the score regression task is selected (and 00 otherwise), and similarly, $\mathds{1}_{\text{deg}}$ equals $1$ if the degradation perception task is selected (and 00 otherwise). After computing rewards for all generated responses $\{r^{(1)},r^{(2)},...,r^{(N)}\}$, the policy model is updated following Eqs. (1) and (3.1).
@@ 3.4 Data Construction
We construct multi-modal training data to jointly train Q-Insight on the score regression and degradation perception tasks. The prompts designed for each task are detailed in Tab. 1. Specifically, a general system prompt is shared across tasks, which encourages the model to explicitly output its reasoning process and provide structured responses. This general prompt is supplemented by task-specific prompts tailored for score regression and degradation perception, respectively. For the score regression task, the input includes a task-specific prompt and the image to be rated, with the Mean Opinion Score (MOS) serving as the ground-truth. In the degradation perception task, the input consists of a prompt and an image characterized by a specific distortion class and severity level. We define five distortion categories: “noise”, “blur”, “JPEG”, “darken”, and “null”, where “null” indicates no distortion. Each distortion type has five severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. The distortion class and corresponding severity level constitute the ground-truth labels. Additionally, for the zero-shot comparative reasoning scenario, inputs include a prompt, two images to be compared, and an optional reference image.
Table 2: PLCC / SRCC comparison on the score regression tasks between our Q-Insight and other competitive IQA methods. All methods except handcrafted ones are trained on the KonIQ dataset. Our Q-Insight outperforms all baseline methods across nearly all benchmarks.

### v2

@@ Abstract
Image quality assessment (IQA) focuses on the perceptual visual quality of images, playing a crucial role in downstream tasks such as image reconstruction, compression, and generation. The rapid advancement of multi-modal large language models (MLLMs) has significantly broadened the scope of IQA, moving toward comprehensive image quality understanding that incorporates content analysis, degradation perception, and comparison reasoning beyond mere numerical scoring. Previous MLLM-based methods typically either generate numerical scores lacking interpretability or heavily rely on supervised fine-tuning (SFT) using large-scale annotated datasets to provide descriptive assessments, limiting their flexibility and applicability. In this paper, we propose Q-Insight, a reinforcement learning-based model built upon group relative policy optimization (GRPO), which demonstrates strong visual reasoning capability for image quality understanding while requiring only a limited amount of rating scores and degradation labels. By jointly optimizing score regression and degradation perception tasks with carefully designed reward functions, our approach effectively exploits their mutual benefits for enhanced performance. Extensive experiments demonstrate that Q-Insight substantially outperforms existing state-of-the-art methods on both score regression and degradation perception tasks, while exhibiting impressive zero-shot generalization and superior comparison reasoning capability. The code and models will be made available.
@@ 1 Introduction
Figure 1: PLCC comparisons between our proposed Q-Insight and existing IQA metrics (left) and three example applications of our Q-Insight (right) are presented. Q-Insight demonstrates significantly improved performance compared to existing methods such as DeQA-Score [58], especially on out-of-domain datasets (e.g., CSIQ [20]). Additionally, Q-Insight effectively supports quality score regression, image degradation perception, and zero-shot image comparison reasoning tasks.
@@
Specifically, we design Q-Insight upon the GRPO framework. In our Q-Insight, we jointly optimize score regression and degradation perception tasks, and carefully design three reward functions: a verifiable score reward for the score regression task, and degradation classification and intensity perception rewards for the degradation perception task. Consequently, Q-Insight effectively exhibits robust reasoning performance using only limited Mean Opinion Scores (MOS) and degradation labels. As shown in Fig. 1, our Q-Insight delivers remarkable performance improvements especially on OOD datasets, while demonstrating comprehensive capabilities across multiple quality assessment and reasoning tasks. For example, it can accurately identify cases where a slightly blurred background, usually regarded as undesirable, effectively helps to emphasize the primary subject of the image. Our empirical investigation reveals that: (1) training solely with score labels results in poor perception of image detail degradations (e.g., JPEG compression), while jointly training with the degradation perception task significantly enhances the model’s sensitivity to such degradations, and (2) score regression and degradation perception tasks are mutually beneficial. Extensive experiments across score regression and degradation perception tasks demonstrate that Q-Insight consistently outperforms existing model-based IQA metrics as well as SFT-driven large language models. Moreover, it exhibits impressive zero-shot generalization on unseen tasks, such as image comparison reasoning, highlighting the robustness and versatility of our method. In summary, our contributions are:
❑ (1) We propose Q-Insight, the first reasoning-style multi-modal large language model specifically designed for comprehensive image quality understanding. Unlike previous methods that depend heavily on detailed textual descriptions for supervised fine-tuning (SFT), our approach achieves superior understanding capability using only limited mean opinion scores or degradation labels.
@@ 2 Related Work
Score-based IQA methods include full-reference and no-reference approaches. Full-reference methods [50, 42, 65] assess image quality by comparing distorted images with high-quality references using traditional metrics (e.g., SSIM [50]) or advanced deep-learning-based metrics [3, 4, 10, 9, 13, 36] like LPIPS [66]. Non-reference methods evaluate quality without reference images, shifting from traditional handcrafted statistics [27, 28, 29, 30, 31, 38] to deep-learning-derived quality priors [18, 19, 24, 33, 44, 72, 73, 45, 49]. Recent multi-modal large language model (MLLM)-based methods, such as Q-Align [53] and DeQA-Score [58], leverage MLLMs’ knowledge and perceptual abilities to produce scores. However, they sacrifice the intrinsic descriptive capabilities of MLLMs.
@@ 3.1 Preliminaries
Group Relative Policy Optimization (GRPO) is an innovative reinforcement learning paradigm that has been widely used in models such as DeepSeek R1-Zero [15], achieving excellent results. Unlike Proximal Policy Optimization (PPO) [39], which requires an explicit critic model to evaluate the performance of the policy model, GRPO [40] directly computes the advantage by comparing a group of responses sampled from the policy model, greatly reducing the computational burden. Specifically, given a query $q$, GRPO samples $N$ distinct responses $\{o^{(1)},o^{(2)},\ldots,o^{(N)}\}$ from the old policy $\pi_{\theta_{\text{old}}}$. Then, the method performs the corresponding actions and receives the respective rewards $\{r^{(1)},r^{(2)},...,r^{(N)}\}$ according to the task-specific rules. By calculating the mean and standard deviation of the rewards, the relative advantages of each response can be obtained as follows:
@@
where $\hat{A}^{(i)}$ represents the normalized relative quality of the $i$-th answer. Overall, GRPO guides the policy model to prioritize higher-quality answers that receive higher reward values within the group. After obtaining the advantage $\hat{A}^{(i)}$, GRPO calculates the ratio of the probabilities of each response under the new policy $\pi_{\theta_{\text{new}}}$ and the old policy $\pi_{\theta_{\text{old}}}$, denoted as $\rho^{(i)}$. To prevent overly large updates to the model and stabilize training, GRPO restricts the $\rho^{(i)}$ to the range $[1-\delta,1+\delta]$. To further maintain closeness to the reference distribution $\pi_{\text{ref}}$, a KL divergence penalty weighted by $\beta$ is adopted. Finally, the optimization objective of GRPO can be formulated as follows:
$\mathcal{J}(\theta)=\mathbb{E}_{[q\sim Q,o^{(i)}\sim\pi_{\theta_{\text{old}}}]}$$\left\{\min\left[\rho^{(i)}\hat{A}^{(i)},\operatorname{clip}\left(\rho^{(i)},1-\delta,1+\delta\right)\hat{A}^{(i)}\right]-\beta\cdot\mathbb{D}_{\text{KL}}[\pi_{\theta_{\text{new}}}||\pi_{\text{ref}}]\right\}$ (2)
where $\rho^{(i)}=\pi_{\theta_{\text{new}}}(o^{(i)}~{}|~{}q)~{}/~{}{\pi_{\theta_{\text{old}}}(o^{(i)}~{}|~{}q)}$, $Q$ denotes the candidate question set, and $\mathbb{D}_{\text{KL}}$ denotes the KL regularization. $\pi_{\text{ref}}$ is typically a frozen pre-trained MLLM. GRPO effectively integrates consistent policy updates and strong reward signals in a balanced way. To our knowledge, we are the first to apply GRPO to image quality understanding tasks, enabling our model to achieve robust reasoning and generalization performance without heavy reliance on extensive annotated data.
@@ 3.2 Overview of Q-Insight
Figure 2: Overview of the proposed Q-Insight framework. The policy model receives queries from multiple tasks and generates corresponding groups of responses accompanied by explicit reasoning steps. Task-specific reward functions ($R_{\text{scr}}$, $R_{\text{deg}}$, and $R_{\text{lev}}$) are then applied, and the policy model is subsequently optimized jointly using the multi-task group relative policy optimization algorithm.
@@ 3.3 Multi-Task Group Relative Policy Optimization
As depicted in Fig. 2, for each input data pair, the policy model $\pi_{\theta}$ generates a group of $N$ responses, denoted as $\{o^{(i)}\}_{i=1}^{N}$. We then evaluate each of these responses using the proposed reward functions ($R_{\text{scr}}$, $R_{\text{deg}}$, and $R_{\text{lev}}$) and obtain the overall rewards $\{r^{(i)}\}_{i=1}^{N}$. Proper reward design is crucial, as informative and carefully constructed rewards directly facilitate Q-Insight’s ability to effectively learn reasoning and perception patterns, thus ensuring robust performance across multiple tasks. Specifically, Q-Insight employs a general format reward function shared across all tasks, as well as task-specific reward functions tailored to the unique characteristics of each individual task.
Format reward evaluates whether the reasoning steps are properly enclosed within “<think>” and “</think>” tags, and the final answer is correctly enclosed within “<answer>” and “</answer>” tags [15]. Additionally, we require that the content inside “<answer>” tags follow a JSON-like format: beginning with “{”, ending with “}”, and containing no additional “{” or “}” characters internally. This ensures that Q-Insight can consistently parse results across different tasks. The reward score $r^{(i)}_{\text{fmt}}$ is set to $1$ if the $i$-th response fulfills all the above conditions; otherwise, its reward is 00.
Rewards for score regression task. A standard way to quantify image quality is by using the Mean Opinion Score (MOS). Instead of directly fitting the MLLM predictions to MOS, we use MOS as a general guideline to motivate the model towards deeper reasoning and generating insightful perspectives during the process of evaluating image quality. Inspired by the treatment of mathematical reasoning tasks in DeepSeek-R1 [15], we consider the continuous MOS prediction as either correct or incorrect, thereby adopting a binary reward to avoid extremely large or small reward values. Denoting the predicted score of the $i$-th response as $\text{scr}_{\text{pred}}^{(i)}$ and the ground-truth score as $\text{scr}_{\text{gt}}$, we design the verifiable reward for scoring as follows. The reward value $r_{\text{scr}}^{(i)}$ for the $i$-th response is determined by:
@@
where $\epsilon$ is a predefined threshold. In particular, if $\epsilon$ is set to 00, the reward simplifies to exact-answer matching. Otherwise, the threshold $\epsilon$ allows the model’s predicted scores to fluctuate within an acceptable range, rather than strictly requiring exact accuracy. As depicted in Fig. 2, the predicted score receives a reward of $1$ if it lies within the threshold $\epsilon$ of the ground-truth MOS, and 00 otherwise.
Rewards for degradation perception task. We find that training solely with score labels leads to poor perception of detailed image degradations (e.g., JPEG compression). This may be because generic multimodal models are pre-trained primarily to capture high-level semantic information, causing them to ignore subtle low-level distortions. To address this issue, we jointly train the model with a degradation perception task, leveraging easily obtainable degradation labels, thus enhancing the model’s sensitivity to these image degradations. In this task, the model is required to predict both the distortion class and the corresponding distortion level. Since distortion class and level are inherently discrete variables, we similarly design binary rewards for this task. Denoting the predicted distortion class and level of the $i$-th response as $\text{deg}^{(i)}_{\text{pred}}$ and $\text{lev}^{(i)}_{\text{pred}}$ respectively, we define the degradation classification reward as follows. The reward value $r^{(i)}_{\text{deg}}$ is determined by:
@@
This means the predicted distortion class receives a reward of $1$ if correct, and 00 otherwise, as illustrated in Fig. 2. Similarly, the intensity perception reward $r^{(i)}_{\text{lev}}$ is determined by:
@@
As depicted in Fig. 2, the predicted distortion level earns a reward of $1$ only when both the predicted class and the level exactly match the ground truth; otherwise, it receives 00.
Overall multi-task reward. Finally, the overall reward of $i$-th response is calculated as:
@@
where $\mathds{1}_{\text{scr}}$ equals $1$ if the score regression task is selected (and 00 otherwise), and similarly, $\mathds{1}_{\text{deg}}$ equals $1$ if the degradation perception task is selected (and 00 otherwise). Note that the reasoning process illustrated in Fig. 2 emerges naturally from the model’s internal capability, without relying on external constraints or additional annotated data. After computing rewards for all generated responses $\{r^{(1)},r^{(2)},...,r^{(N)}\}$, the policy model is updated following Eqs. (1) and (2). With this flexible design, Q-Insight can seamlessly switch between tasks and jointly optimize them during training. During inference, the trained policy model can directly perform image quality understanding without requiring additional fine-tuning. Experimental results presented in Tabs. 4.2 and 4.2 and Fig. A further demonstrate that jointly addressing the score regression and degradation perception tasks substantially improves performance, highlighting the beneficial interactions between these two tasks.
@@ 3.4 Data Construction
We construct multi-modal training data to jointly train Q-Insight on the score regression and degradation perception tasks. The prompts designed for each task are detailed in Tab. A in the appendix. For the score regression task, the input includes a task-specific prompt and the image to be rated, with the Mean Opinion Score (MOS) serving as the guideline to calculate the corresponding reward. In the degradation perception task, the input consists of a prompt and an image characterized by a specific distortion class and severity level. There are five distortion categories: “noise”, “blur”, “JPEG”, “darken”, and “null”, where “null” indicates no distortion. Each distortion type has five severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. The distortion class and corresponding severity level constitute the ground-truth labels to calculate the degradation classification and intensity perception rewards. Overall, our carefully-designed GRPO-based framework and multi-task training strategy help Q-Insight achieve robust reasoning and perception capability even from limited annotated labels. More importantly, this flexible approach facilitates effective generalization to various low-level vision applications, as demonstrated by extensive experiments in Sec. 4, clearly highlighting the advantages of our framework in addressing practical vision tasks.
Table 1: PLCC / SRCC comparison on the score regression tasks between our Q-Insight and other competitive IQA methods. All methods except handcrafted ones are trained on the KonIQ dataset. The best and second-best results of each test setting are highlighted in bold red and underlined blue.